In such an experiment, a brass wire of diameter 9.50 x 10^-4 m is used. If the young modulus for brass is 9.86 x 10^10 Nm^-2, find the greatest force which could be use to keep within the limitation.
1
Expert's answer
2020-08-05T16:42:51-0400

Since Stress "\propto" strain

Then

"Stress = Y(strain)"

"\frac{F}{A} = Y(strain)" where F is the force and A is the area.


then

"F = AY(strain)"

Let strain be 0.0001 for sake of simplicity

then

"F = 9.86*10^{10} * \pi*(\frac{9.5*10^{-4}}{2})^2*0.0001 = 6.98 N"